Factorise the following expressions.
a^2 + 8a + 16 ?????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

a^2 + 8a + 16
= a^2+2×4×a+4^2
= (a+4)^2
Using identity: (x+y)^2= x^2+2xy+y^2

Thanks